# Estimation of Concept Explanations Should be Uncertainty Aware

### 13 Dec 2023 v1

@@ 7 Related Work
Concept Bottleneck Models use a set of predefined human-interpretable concepts as an intermediate feature representation to make the predictions  (Koh et al., 2020; Bau et al., 2017a; Kim et al., 2018; Zhou et al., 2018). CBM allows human test-time intervention which has been shown to improve overall accuracy  (Barker et al., 2023). Traditionally, they require labelled data with concept annotations and typically the accuracy is worse than the standard models without concept bottleneck. To address the limitation of concept annotation, recent works have leveraged large pretrained multi-modal models like CLIP  (Oikarinen et al., 2023; Yuksekgonul et al., 2022). There have also been efforts to enhance the reliability of CBMs by focusing on the information leakage problem  (Havasi et al., 2022; Marconato et al., 2022), where the linear model weights estimated from concept activations utilize the unintended information, affecting the interpretability. Concept Embedding Models (CEM)  (Espinosa Zarlenga et al., 2022) overcome the trade-off between accuracy and interpretability by learning high-dimensional concept embeddings. However, addressing the noise in the concept prediction remains underexplored. Collins et al. (2023) have studied human uncertainty in concept-based models and have shown the importance of considering uncertainty over concepts in improving the reliability of the model. Kim et al. (2023a) proposed the Probabilistic Concept Bottleneck Models (ProbCBM) and is closely related to our work. They too argue for the need to model uncertainty in concept prediction for reliable explanations. However, their method of noise estimation in concept activations requires retraining the model and cannot be applied directly when concept activations are estimated using CLIP. Moreover, they use simple MC sampling to account for noise in concept activations.
@@
Concept based explanations use a separate probe dataset to first learn the concept and then explain through decomposition either the individual predictions or overall label features. Yeh et al. (2022) contains a brief summary of existing concept based explanation methods. Our proposed method is very similar to concept based explanations (CBE) (Kim et al., 2018; Bau et al., 2017a; Zhou et al., 2018; Ghorbani et al., 2019). Ramaswamy et al. (2022a) emphasized that the concepts learned are sensitive to the probe dataset used and therefore pose problems when transferring to applications that have distribution shift from the probe dataset. Moreover, they also highlight other drawbacks of existing CBE methods in that concepts can sometimes be harder to learn than the label itself (meaning the explanations may not be causal) and that the typical number of concepts used for explanations far exceed what a typical human can parse easily. Achtibat et al. (2022) championed an explanation method that provides explanation highlighting important feature (answering “where”) and what concepts are used for prediction thereby combining the strengths of global and local explanation methods. Choi et al. (2023) have built upon the current developments in CBE methods for providing explanations for out-of-distribution detectors. Wu et al. (2023) introduced the causal concept based explanation method (Causal Proxy Model), that provides explanations for NLP models using counterfactual texts. Moayeri et al. (2023) also used CLIP to interpret the representations of a different model trained on uni-modal data.
@@ 8 Conclusion
We studied concept explanation methods with a focus on data-efficient systems that exploit pretrained multi-modal models. We demonstrated with simple examples the reliability challenge of existing estimators of concept explanations and motivated the need for modeling uncertainty in estimation and informing user the uncertainty in importance scores. Accordingly, we proposed an uncertainty-aware and data-efficient estimator called U-ACE, which readily yielded several benefits. We demonstrated the merits of our estimator through theoretical analysis, controlled study experiments and two challenging real-world evaluation with around 700 concepts. To the best of our knowledge, previous evaluations did not consider concept explanations with as many concepts. Our results showed that concept explanations estimated by U-ACE are more reliable.
Limitations and Future Work $\bullet$ The need and advantage when modeling uncertainty is also applicable when learning concept activations using datasets with concept annotations. However, our experimental setup is only focused on using CLIP for specifying concepts. $\bullet$ We did not model the uncertainty in CLIP’s knowledge of a concept. Epistemic uncertainty due to CLIP when modelled may improve reliability further, which we leave for future work.
Appendix
@@ A.1 Derivation of posterior on weights
The result of posterior distribution of weights follows directly from the form of posterior under normal prior on weights as explained as Salakhutdinov (2011) (Slide 10). For the sake of completeness, we also derive the result below.
@@
A.2 Corollary of Proposition 2
Corollary 2.
For the data setup of Proposition 2, the following results holds when $u_{1}=\mathbf{w},\sigma_{1}\approx 0$ and $u_{k}^{T}\mathbf{w}=0,\quad\forall k\in[2,K]$. Then the probability that the standard estimator returns the first concept as the most salient decreases exponentially with the number of concepts. On the other hand, the importance score assigned by U-ACE is 1 for the only relevant first concept and 0 otherwise.
Proof.
Plugging in the values for the special case of $u_{1}=w,\sigma_{1}\approx 0$ and $u_{k}^{T}w=0,k\geq 2$ in the closed form solution from Proposition 2, we have the following results for the standard linear estimator and U-ACE.
Solution of standard linear estimator: $v_{1}=1$ and $v_{k}\sim\mathcal{N}(0,\sigma_{k}^{2}\frac{\|w\|}{\|u_{k}\|^{2}})$ for k2absent2$\geq$ 2≥ 2.
[…]
Solution of U-ACE: $v_{1}=1,v_{2},v_{3},\dots=0$ follows directly from plugging in the values in to result of the proposition.
A.3 Algorithm
@@ Appendix B Maximum Likelihood Estimation of U-ACE parameters
We implement the objective using Pyro software library (Bingham et al., 2019) and Adam optimizer.
@@ Appendix C Proof of Proposition 1
Corresponding to $v_{k}$ in $f$, there must be an equivalent vector $w$ in the embedding space of g.
@@ Appendix D Proof of Proposition 2
The concept importance estimated by U-ACE when the input dimension is sufficiently large and for some $\lambda>0$ is approximately given by $v_{k}=\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{i}^{T}\mathbf{u}_{k}+\lambda\sigma_{k}^{2}}$. On the other hand, the importance scores estimated using vanilla linear estimator under the same conditions is distributed as $v_{k}\sim\mathcal{N}(\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{k}^{T}\mathbf{u}_{k}},\sigma_{k}^{2}\frac{\|w\|^{2}}{\|u_{k}\|^{2}})$.
@@
We use the known result that inner product of two random vectors is close to 0 when the number of dimensions is large, i.e. $u_{i}^{T}u_{j}\approx 0,i\neq j$.
Result with vanilla estimator. We first show the solution using vanilla estimator is distributed as given by the result above. We wish to estimate $v_{1},v_{2},\dots$ such that we approximate the prediction of model-to-be-explained: $y=w^{T}\mathbf{x}$. We denote by $w_{k}$ sampled from the normal distributin of concept vectors. We require $w^{T}\mathbf{x}\approx\sum_{k}v_{k}w_{k}^{T}\mathbf{x}$. In effect, we are optimising for $v$s such that $\|w-\sum_{k}v_{k}w_{k}\|^{2}$ is minimized. We multiply the objective by $u_{k}$ and use the result that random vectors are almost orthogonal in high-dimensions to arrive at objective $\operatorname*{arg\,min}_{v_{k}}\|w_{k}^{T}w-v_{k}(w_{k}^{T}w_{k})\|$. Which is minimized trivially when $v_{k}=\frac{w_{k}^{T}w}{\|w_{k}\|^{2}}$. Since $w_{k}$ is normally distributed with $\mathcal{N}(u_{k},\sigma_{k}^{2}I)$, $w_{k}^{T}w=(u_{k}+\epsilon)^{T}w,\quad\epsilon\sim\mathcal{N}(0,I)$ is also normally distributed with $\mathcal{N}(u_{k}^{T}w,\sigma_{k}^{2}\|w\|^{2})$. We approximate the denominator with its average and ignoring its variance, i.e. $\|w_{k}\|^{2}=\mathcal{N}(\|u_{k}\|^{2},\sigma_{k}^{2})\approx\|u_{k}\|^{2}$ which is when $\|u_{k}\|^{2}>>\sigma^{2}$. We therefore have the result on distribution of $v_{k}$.
Using U-ACE. Similar to vanilla estimator, U-ACE optimizes $v_{k}$ using the following objective.
$\ell=\operatorname*{arg\,min}_{v}\{\|w-\sum_{k}v_{k}u_{k}\|^{2}+\lambda\sum_{k}\sigma_{k}^{2}v_{k}^{2}\}$
@@
$-u_{k}^{T}(w-\sum_{j}v_{j}u_{j})+\lambda\sigma_{k}^{2}v_{k}=0$
$\implies v_{k}=\frac{u_{k}^{T}w}{\|u_{k}\|^{2}+\lambda\sigma_{k}^{2}}$
Appendix E Proof of Proposition 3
The importance score, denoted $v_{1},v_{2}$, estimated by U-ACE are bounded from above by $\frac{1}{N\lambda}$, i.e. $v_{1},v_{2}=\mathcal{O}(1/N\lambda)$ where $\lambda>0$ is a regularizing hyperparameter and N the number of examples.
@@ Appendix E Proof of Proposition 3
We first show that the values of $v_{1},v_{2}$ in closed form are as below before we derive the final result.
$v_{1}$ $=\frac{\frac{S_{1}}{S_{2}}(1-\beta_{2})^{2}}{\frac{S_{1}}{S_{2}}(\beta_{2}^{2}(1-\beta_{1})^{2}+\beta_{1}^{2}(1-\beta_{2})^{2})+\lambda(1-\beta_{1})(1-\beta_{2})}$
$v_{2}$ $=\frac{\frac{S_{1}}{S_{2}}(1-\beta_{1})^{2}}{\frac{S_{1}}{S_{2}}(\beta_{1}^{2}(1-\beta_{2})^{2}+\beta_{2}^{2}(1-\beta_{1})^{2})+\lambda(1-\beta_{1})(1-\beta_{2})}$
where $S_{1}=\sum_{i}y_{1}$, $S_{2}=\sum_{i}y_{i}^{2}$ and $\lambda>0$ is a regularizing hyperparameter.
We then observe that if $\mathbf{x}$ is normally distributed then $y=w^{T}\mathbf{x}$ is also normally distributed with the value of $\frac{S_{1}}{S_{2}}$ is of the order $\mathcal{O}(1/N)$. Since $\beta_{1},\beta_{2}$ are very close to 0, we can approximate the expression for $v_{1}$ as below.
$v_{1}$ $\approx\frac{S_{1}}{S_{2}}(1-\beta_{2})^{2}\frac{1}{\lambda(1-\beta_{1})(1-\beta_{2})}=\mathcal{O}(1/N\lambda)$
Importance scores from a standard estimator.
$\text{When }c_{1}^{(1)}=(\beta_{1}u+(1-\beta_{1})v)^{T}z^{(i)},\quad c_{2}^{(i)}=(\beta_{2}u+(1-\beta_{2})v)^{T}z^{(i)}$
@@
$\frac{(1-\beta_{2})c_{1}-(1-\beta_{1})c_{2}}{(1-\beta_{2})\beta_{1}-(1-\beta_{1})\beta_{2}}=\frac{(1-\beta_{2})c_{1}-(1-\beta_{1})c_{2}}{\beta_{1}-\beta_{2}}=u^{T}z_{i}=y_{i}$
$\implies\frac{1-\beta_{2}}{\beta_{1}-\beta_{2}}c_{1}+\frac{1-\beta_{1}}{\beta_{1}-\beta_{2}}c_{2}=y_{i}$
@@
Appendix F Additional experiment details
List of fruit concepts from Section 4.1.

### 05 Apr 2024 v2

@@ 7 Related Work
Concept Bottleneck Models use a set of predefined human-interpretable concepts as an intermediate feature representation to make the predictions  [Koh et al., 2020, Bau et al., 2017a, Kim et al., 2018, Zhou et al., 2018]. CBM allows human test-time intervention which has been shown to improve overall accuracy  [Barker et al., 2023]. Traditionally, they require labelled data with concept annotations and accuracy is typically worse than the standard models without concept bottleneck. To address the limitation of concept annotation, recent works have leveraged large pretrained multimodal models like CLIP  [Oikarinen et al., 2023, Yuksekgonul et al., 2022]. Concept Embedding Models (CEM)  [Espinosa Zarlenga et al., 2022] overcome the trade-off between accuracy and interpretability by learning high-dimensional concept embeddings. However, addressing the noise in the concept prediction remains underexplored. Collins et al. [2023] have studied human uncertainty in concept-based models and elucidate the importance of considering uncertainty over concepts in improving the reliability of the model. Closely related to our work, Kim et al. [2023a] propose the Probabilistic Concept Bottleneck Models (ProbCBM). They too argue for the need to model uncertainty in concept prediction for reliable explanations. However, their method of noise estimation in concept activations requires retraining the model and cannot be applied directly when concept activations are estimated using CLIP. Moreover, they use simple MC sampling to account for noise in concept activations, which is not nearly as effective (Appendix H.1).
@@
Concept-based explanations use a separate probe dataset to first learn the concept and then explain through decomposition either the individual predictions or overall label features. Yeh et al. [2022] contains a brief summary of existing concept based explanation methods. Our proposed method is very similar to concept-based explanations (CBE) [Kim et al., 2018, Bau et al., 2017a, Zhou et al., 2018, Ghorbani et al., 2019]. Ramaswamy et al. [2022a] emphasised that the concepts learned are sensitive to the probe dataset used and therefore pose problems when transferring to applications that have distribution shift from the probe dataset. They further highlight the drawback of existing CBE methods that concepts can sometimes be harder to learn than the label itself (meaning the explanations may not be causal) and that the typical number of concepts used for explanations far exceed what a typical human can parse easily. Achtibat et al. [2022] championed an explanation method that provides explanation highlighting important feature (answering “where”) and what concepts are used for prediction thereby combining the strengths of global and local explanation methods. Choi et al. [2023] have built upon the current developments in CBE methods for providing explanations for out-of-distribution detectors. Wu et al. [2023] introduced the causal concept based explanation method (Causal Proxy Model), that provides explanations for NLP models using counterfactual texts. Moayeri et al. [2023] also used CLIP to interpret the representations of a different model trained on uni-modal data.
@@ 8 Conclusion
We studied concept explanation methods with a focus on data-efficient systems that exploit pretrained multimodal models. We highlighted the quality challenge of the of existing estimators of concept explanations via simple examples and motivated the need for modelling uncertainty in their estimation. Accordingly, we proposed an uncertainty-aware and data-efficient estimator called U-ACE. We demonstrated the merits of our estimator through theoretical analysis, controlled study experiments and three challenging real-world evaluation with order of thousand concepts. Our results establish the strong promise of concept explanations estimated using our Bayesian method for effective model debugging.
Limitations and Future Work. Pretrained multimodal models enabled us to work with an open-world specification of the set of concepts, without necessitating expensive annotation data. Yet, dependence on a pretrained model may limit application in certain specialised domains, which while emerging in domains like healthcare  [Huang et al., 2023], may temporarily hinder widespread adoption. In the same vein, an estimator that also models the uncertainty due to lack of knowledge of the pretrained multimodal model about a concept will be better suited for addressing any lapses in the pretrained model, which we leave for the future. Additionally, while we do not run user studies in this work, a natural and exciting next step is to explore whether U-ACE explanations are preferred by real humans.
@@ A.1 Derivation of posterior on weights
The result of posterior distribution of weights follows directly from the form of posterior under normal prior on weights as explained as Salakhutdinov [2011] (Slide 10). For the sake of completeness, we also derive the result below.
@@
A.2 Algorithm
We describe the algorithm summarizing U-ACE in 1. An additional technical detail of the algorithm is a step to sparsify weights as described below.
Sparsifying weights for interpretability. As a dense weight matrix can be hard to interpret, we induce sparsity in $W_{c}$ by setting all the values below a threshold to zero. We pick the threshold such that the accuracy on train split does not fall by more than $\kappa$, which is a positive hyperparameter.
@@ Appendix B Maximum Likelihood Estimation of U-ACE parameters
We implement the objective using Pyro software library [Bingham et al., 2019] and Adam optimizer.
@@ Appendix C Proof of Proposition 1
[…]
@@ Appendix D Proof of Proposition 2
The concept importance estimated by U-ACE when the input dimension is sufficiently large and for some $\lambda>0$ is approximately given by $v_{k}=\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{k}^{T}\mathbf{u}_{k}+\lambda\sigma_{k}^{2}}$. On the other hand, the importance scores estimated using vanilla linear estimator under the same conditions is distributed as $v_{k}\sim\mathcal{N}(\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\mathbf{u}_{k}^{T}\mathbf{u}_{k}},\sigma_{k}^{2}\frac{\|\mathbf{w}\|^{2}}{\|\mathbf{u}_{k}\|^{2}})$.
@@
We use the known result that inner product of two random vectors is close to 0 when the number of dimensions is large, i.e. $\mathbf{u}_{i}^{T}\mathbf{u}_{j}\approx 0,i\neq j$.
Solution with standard estimator. We first show the solution using vanilla estimator is distributed as given by the result above. We wish to estimate $v_{1},v_{2},\dots$ such that we approximate the prediction of model-to-be-explained: $y=\mathbf{w}^{T}\mathbf{x}$. We denote by $\mathbf{w}_{k}$ sampled from the normal distribution of concept vectors. We require $w^{T}\mathbf{x}\approx\sum_{k}v_{k}\mathbf{w}_{k}^{T}\mathbf{x}$. In effect, we are optimising for $v$s such that $\|\mathbf{w}-\sum_{k}v_{k}\mathbf{w}_{k}\|^{2}$ is minimized. We multiply the objective by $\mathbf{u}_{k}$ and use the result that random vectors are almost orthogonal in high-dimensions to arrive at objective $\operatorname*{arg\,min}_{v_{k}}\|\mathbf{w}_{k}^{T}\mathbf{w}-v_{k}(\mathbf{w}_{k}^{T}\mathbf{w}_{k})\|$. Which is minimized trivially when $v_{k}=\frac{\mathbf{w}_{k}^{T}\mathbf{w}}{\|\mathbf{w}_{k}\|^{2}}$. Since $\mathbf{w}_{k}$ is normally distributed with $\mathcal{N}(\mathbf{u}_{k},\sigma_{k}^{2}I)$, $\mathbf{w}_{k}^{T}\mathbf{w}=(\mathbf{u}_{k}+\vec{\epsilon})^{T}\mathbf{w},\quad\vec{\epsilon}\sim\mathcal{N}(0,I)$ is also normally distributed with $\mathcal{N}(\mathbf{u}_{k}^{T}\mathbf{w},\sigma_{k}^{2}\|\mathbf{w}\|^{2})$. We approximate the denominator with its average and ignoring its variance, i.e. $\|\mathbf{w}_{k}\|^{2}=\mathcal{N}(\|\mathbf{u}_{k}\|^{2},\sigma_{k}^{2})\approx\|\mathbf{u}_{k}\|^{2}$ which is when $\|\mathbf{u}_{k}\|^{2}>>\sigma^{2}$. We therefore have the result that $v_{k}$ is normally distributed with mean $\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\|\mathbf{u}_{k}\|^{2}}$ and variance $\sigma_{k}^{2}\frac{\|\mathbf{w}\|^{2}}{\|\mathbf{u}_{k}\|^{2}}$.
Solution with U-ACE. Unlike the standard estimator, U-ACE seeks a solution by first estimating the distribution on concept activations. We make an approximation of our proposed estimation of U-ACE and find a solution using the following objective, recall that the $k^{th}$ activation vector is defined to have been sampled from $\mathcal{N}(\mathbf{u}_{k},\sigma_{k}^{2}I)$.
$\ell=\operatorname*{arg\,min}_{v}\{\|\mathbf{w}-\sum_{k}v_{k}\mathbf{u}_{k}\|^{2}+\lambda\sum_{k}\sigma_{k}^{2}v_{k}^{2}\}$
With sufficient number of examples, the mean and variance $\mathbf{u}_{k},\sigma^{2}$ can be faithfully estimated. The above objective captures the essence of U-ACE(Eqn. (1)) because (1) it fits the observations, and (2) regularizes importance of concepts with high variance. We now proceed with finding the closed form solution of $v_{k}$.
$\ell=\operatorname*{arg\,min}_{v}\{\|\mathbf{w}-\sum_{k}v_{k}\mathbf{u}_{k}\|^{2}+\lambda\sum_{k}\sigma_{k}^{2}v_{k}^{2}\}$
setting $\text{setting }\frac{\partial\ell}{\partial v_{k}}=0\text{ we obtain}$
$-\mathbf{u}_{k}^{T}(w-\sum_{j}v_{j}\mathbf{u}_{j})+\lambda\sigma_{k}^{2}v_{k}=0$
and using almost zero inner product result stated above, i.e. $\mathbf{u}_{i}^{T}\mathbf{u}_{j}\approx 0,i\neq j$, we get
$\implies v_{k}=\frac{\mathbf{u}_{k}^{T}\mathbf{w}}{\|\mathbf{u}_{k}\|^{2}+\lambda\sigma_{k}^{2}}$
Appendix E Corollary of Proposition 2
Corollary 2.
For the data setup of Proposition 2, the following results holds when $\mathbf{u}_{1}=\mathbf{w},\sigma_{1}\approx 0$ and $\mathbf{u}_{k}^{T}\mathbf{w}=0,\quad\forall k\in[2,K]$. Then the probability that the standard estimator returns the first concept as the most salient decreases exponentially with the number of concepts. On the other hand, the importance score assigned by U-ACE is 1 for the only relevant first concept and 0 otherwise.
Proof.
Plugging in the values for the special case of $\mathbf{u}_{1}=\mathbf{w},\sigma_{1}\approx 0$ and $\mathbf{u}_{k}^{T}\mathbf{w}=0,k\geq 2$ in the closed form solution from Proposition 2, we have the following results for the standard linear estimator and U-ACE.
Solution of standard estimator. $v_{1}=1$ and $v_{k}\sim\mathcal{N}(0,\sigma_{k}^{2}\frac{\|\mathbf{w}\|^{2}}{\|\mathbf{u}_{k}\|^{2}})$ for k2absent2$\geq$ 2≥ 2.
@@ Appendix E Corollary of Proposition 2
Solution of U-ACE. $v_{1}=1,v_{2},v_{3},\dots=0$ follows directly from plugging in the values in to result of the proposition.
Appendix F Proof of Proposition 3
The importance score estimated by U-ACE is approximately $\frac{b_{1}/2}{1+\lambda\sigma^{2}},\frac{b_{2}/2}{1+\lambda\sigma^{2}}$, where $\lambda>0$ is a regularizing hyperparameter.
@@ Appendix F Proof of Proposition 3
Since $\beta_{1},\beta_{2}$ are normal distributed, the concept activation vectors $\beta_{1}\mathbf{u}+(1-\beta_{1})\mathbf{v}$ and $\beta_{2}\mathbf{u}+(1-\beta_{2})\mathbf{v}$ are also normally distributed. We derive their closed form below. Recall that $\beta_{1}\sim\mathcal{N}(b_{1},\sigma^{2}),\beta_{2}\sim\mathcal{N}(b_{2},\sigma^{2})$ and that $\mathbf{u},\mathbf{v}$ are orthogonal and unit vectors.
$\mathbb{E}[\beta_{1}\mathbf{u}+(1-\beta_{1})\mathbf{v}]$ $=b_{1}\mathbf{u}+(1-b_{1})\mathbf{v}$
$\mathbb{V}[\beta_{1}\mathbf{u}+(1-\beta_{1})\mathbf{v}]$ $=\mathbb{E}[(\beta_{1}-b_{1})\mathbf{u}+(b_{1}-\beta_{2})\mathbf{v})^{2}]$
$=\mathbb{E}[(\beta_{1}-b_{1})^{2}]+\mathbb{E}[(\beta_{2}-b_{2})^{2}]$
$=2\sigma^{2}$
As argued of the objective to approximate U-ACE in Appendix D, we may obtain the closed form solution for the concept scores $\eta_{1},\eta_{2}$ by solving the following objective.
$\ell=\|\mathbf{u}-\eta_{1}(b_{1}\mathbf{u}+(1-b_{1})\mathbf{v})-\eta_{2}(b_{2}\mathbf{u}+(1-b_{2})\mathbf{v})\|^{2}+2\lambda\sigma^{2}(\eta_{1}^{2}+\eta_{2}^{2})$
By setting $\partial\ell/\partial\eta_{1}=0$, $\partial\ell/\partial\eta_{2}=0$, and simplifying, we get the following.
$b_{1}$ $=\eta_{1}(b_{1}^{2}+(1-b_{1})^{2}+2\lambda\sigma^{2})+\eta_{2}(b_{1}b_{2}+(1-b_{1})(1-b_{2}))$
$b_{2}$ $=\eta_{2}(b_{2}^{2}+(1-b_{2})^{2}+2\lambda\sigma^{2})+\eta_{1}(b_{1}b_{2}+(1-b_{1})(1-b_{2}))$
Eliminating variables, and solving for $\eta_{1}$, we get
$\eta_{1}$ $=\frac{2\lambda b_{1}\sigma^{2}+(1-b_{2})(b_{1}-b_{2})}{(b_{1}-b_{2})(b_{1}+b_{2}-2b_{1}b_{2})+2\lambda\sigma^{2}(2\lambda\sigma^{2}+b_{1}^{2}+b_{2}^{2}+(1-b_{1})^{2}+(1-b_{2})^{2})}$ $=\text{similarly for $\eta_{2}$}$
By substituting $b_{1}-b_{2}\approx 0$, the expressin can be simplified as
$\eta_{1}$ $=\frac{b_{1}}{2\lambda\sigma^{2}+b_{1}^{2}+b_{2}^{2}+(1-b_{1})^{2}+(1-b_{2})^{2}}$
If we now use the assumption that $b_{1},b_{2}\approx 0$, $\therefore 1-b_{1},1-b_{2}\approx 1$, we get the final form below.
$\eta_{1}$ $=\frac{b_{1}}{2(1+\lambda\sigma^{2})}$
$\eta_{2}$ $=\frac{b_{2}}{2(1+\lambda\sigma^{2})}$
Solution of standard estimator.
$\text{When }c_{1}^{(i)}=(\beta_{1}u+(1-\beta_{1})v)^{T}z^{(i)},\quad c_{2}^{(i)}=(\beta_{2}u+(1-\beta_{2})v)^{T}z^{(i)}$
@@
$\frac{(1-\beta_{2})c_{1}^{(i)}-(1-\beta_{1})c_{2}^{(i)}}{(1-\beta_{2})\beta_{1}-(1-\beta_{1})\beta_{2}}=\frac{(1-\beta_{2})c_{1}^{(i)}-(1-\beta_{1})c_{2}^{(i)}}{\beta_{1}-\beta_{2}}=u^{T}z^{(i)}=y^{(i)}$
$\implies\frac{1-\beta_{2}}{\beta_{1}-\beta_{2}}c_{1}^{(i)}+\frac{1-\beta_{1}}{\beta_{1}-\beta_{2}}c_{2}^{(i)}=y_{i}$
@@
Appendix G Additional experiment details
G.1 Simulated Study
In this section, we consider explaining a two-layer CNN model trained to classify between solid color images with pixel noise as shown in Figure 4. The colors on the left (red; green) are defined as label 0 and the colors on the right (blue; white) are defined as label 1.
Figure 4: Toy
The model-to-be-explained is trained on a dataset with equal proportion of all colors; we therefore expect that all constituent colors of a label are equally important for the label. We specify a concept set with the four colors encoded by their literal name red, green, blue, white. U-ACE (along with others) attribute positive importance for red, green and negative or zero importance for blue, white when explaining label 0 using a concept set with only the four task-relevant concepts and when the probe-dataset is the same distribution as the the training dataset. However, quality of explanations quickly degrades when the probe-dataset is shifted, or if the concept set is misspecified.
Unreliability due to dataset shift.
We varied the probe-dataset to include varying populations of colors while keeping the concept set and model-to-be-explained fixed. We observed that importance of a concept estimated with standard CBEs varied with the choice of probe-dataset for the same underlying model-to-be-explained as shown in left and middle plots of Figure 5. Most methods attributed incorrect importance to the red concept when it is missing (left extreme of left plot), and similarly for the green concept (left extreme of middle plot). Such explanations would have led the user to believe that green is more important than red, or red is more important than green, depending on the probe-dataset used (as shown in the right most plot). In practice, mis-communicated concept importance could be disastrous, if such importances are used to inform decision making and/or discovery. Because U-ACE also informs the user of uncertainty in the estimated importance, we see that the difference in importance scores between the two colors at either extremes is not statistically significant as shown in the rightmost plot.
Figure 6: U-ACE is reliable even with overly complete concept set.
Over-complete concept set. We now evaluate the quality of explanations when the concept set is misspecified. We induce mispecification by making the concept set over-complete by gradually expanding it to include common fruit names (Appendix G contains the full list), which are clearly irrelevant to the task. We obtain the explanations using an in-distribution probe-dataset that contains all colors in equal proportion. Figure 6 shows the score of most salient fruit concept with increasing number of fruit (nuisance) concepts on X-axis. We observe that U-ACE is far more robust to the presence of nuisance concepts. Robustness to irrelevant concepts is important because it allows the user to begin with a superfluous set of concepts and find their relevance to model-to-be-explained, instead of requiring users to guess relevant concepts, which is ironically the very purpose of using concept explanations.
Concept Y-CBM O-CBM U-ACE
red or blue -75.4 -1.8 0.1
blue or red 21.9 -1.9 0
green or blue -1.4 1.6 0
blue or green -23.1 1.6 0
List of fruit concepts from Section G.1.